find the product shown below in simplest form:
(x+5) (x-5)
explain how you can use this product to determine that (15) (5)= 75

Found 2 solutions by Fombitz, MathTherapy:
Answer by Fombitz(32388)
Answer by MathTherapy(10555)


find the product shown below in simplest form:
(x+5) (x-5)
explain how you can use this product to determine that (15) (5)= 75
let x = 10
Then (x + 5)(x - 6) becomes: (10 + 5)(10 - 5), or: (15)(5) = 75
Also could be calculated as: 10(10 - 5) + 5(10 - 5) = 100 - 50 + 50 - 25 = 75